Carbon dioxide gas (CO2) enters a compressor with a flowrate of 0.1 kg/s at 5ºC, 100 kPa and exits at 125ºC, 500 kPa. Determine the rate of heat loss to the surroundings, in kW, if the compressor requires 12 kW of power to operate. (You will need to be careful to avoid a sign error. Remember that by our sign convention cooling is negative and that power into a system is also negative.)
Answer #1

Q. Given, Pi= lookpa wer) = 12kW Ti = 5C=23317 Compressor LG Resookpa Te = 125°C= 3966 mco.ilgis Erit Entry Treating coz gasCoa VR : 1.33 (0.18895) = 0.7615 KJlg.k 1-1 1-33-1 0.1(0.7615)(278) + 8 = 0.100.7615)(398) – 12 á= -2.862 KW -ve sign indicat